OpenStudy (anonymous):

Joan’s washing machine tub is a cylinder with a height of 17 in. and a diameter of 22 in. Her washer was half full of water when it stopped working. The repair person’s pump removed 46.2 in3 of water every second. How many seconds did it take to empty the washer? Use 3.14 to approximate pi, and round your answer to the nearest whole number.

OpenStudy (anonymous):

Use the volume for a cylinder V=pi*r^2*h. Divide by 2 since it is half full. Then take that answer and divide it by the amount of water the pump removes per second

OpenStudy (anonymous):

don't forget r= diameter/2

OpenStudy (anonymous):

598.35

OpenStudy (jhannybean):

If volume of a cylinder is defined as: \(V = \pi r^2 h\), Because they've given you diameter, you want to find the radius as the diameter / 2 : \(r = \dfrac{D}{2}\) then removing water has to be with the time, you want to divide how much volume in contained within the washer itself up to the time it's stopped working, which would be \(V_{\sf half ~full} = \pi r^2 h_1\) You're given \(h=17 ~in~,~ r=\dfrac{D}{2} = \dfrac{22~ in}{2} = 11~ in\) , Adn so you can use this to find \(V_{\sf half~ full}\) Then to find out how much TIME, make a ratio: \(\dfrac{V_{\sf half~full}}{V_{\sf rate}}\)_

OpenStudy (jhannybean):

\[\frac{V_{\sf half~full} ~in^3}{V_{\sf rate} ~\dfrac{in^3}{s}} = ~? ~ s \]

OpenStudy (jhannybean):

Nope.

OpenStudy (jhannybean):

Find your \(V_{\sf half~full}\) and tell me what you get using \(\pi = 3.14\).

OpenStudy (anonymous):

what's the Vhalf full

OpenStudy (anonymous):

Vhalf full should be volume of the entire washing machine divided by two

OpenStudy (anonymous):

soooooooooo

OpenStudy (anonymous):

93.5

OpenStudy (anonymous):

V=(3.14)*(11)^2*(17)=?

OpenStudy (anonymous):

3485.98

OpenStudy (anonymous):

6485.98 sorry type-o

OpenStudy (anonymous):

Great that is that volume, since it is half full you divide that by 2

OpenStudy (anonymous):

Now you have 3229.49 in^3 * 1/46.2 seconds/in^3 = The inches should cancel out and you are left with your answer in seconds

OpenStudy (jhannybean):

Precisely.

OpenStudy (anonymous):

i don't know the answer

OpenStudy (anonymous):

can't figure it out

OpenStudy (anonymous):

just tell me the answer

OpenStudy (jhannybean):

Sorry,I cant do that.

OpenStudy (anonymous):

it only lets me type 2 numbers

OpenStudy (anonymous):

If you read the problem its says round to the nearest second

OpenStudy (anonymous):

69

OpenStudy (anonymous):

or 70

OpenStudy (jhannybean):

Solve: \[\sf 3229.49~~ in^3 \cdot \frac{1~s}{46.2 ~in^3}\]

OpenStudy (anonymous):

What is the rule for rounding. 5-9 round up and for 0-4 round down.

OpenStudy (anonymous):

69.91 rounded up is 70

OpenStudy (jhannybean):

Yep.

OpenStudy (anonymous):

70 is the answer, really

OpenStudy (anonymous):

u both were wrong! it was 7

OpenStudy (anonymous):

70